# Large Language Models and Code Security: A Systematic Literature Review

### v2

@@ Abstract
I Introduction
Large Language Models (LLMs) are Machine Learning (ML) models designed for natural language processing [1]. Representative models such as ChatGPT [2], Llama [3], GitHub Copilot [4], and BERT [5] have gained extreme popularity and recognition in the last couple of years due to their ability to perform well across different types of tasks.
Among other capabilities, LLMs are effective in translating natural language explanations into code, assisting in code comprehension, debugging, and answering questions related to code [6]. Many developers utilize these abilities to optimize their coding processes, reducing development time and thereby improving productivity. According to GitHub statistics, GitHub Copilot is now responsible for generating approximately 46% of code and boosts developers’ coding speed by up to 55% [7]. In addition to their coding skills, LLMs can be powerful tools for vulnerability detection and fixing, particularly when provided with detailed prompts and clear background information [8][9].
While LLMs represent an advancement in code generation, they exhibit several limitations that pose security risks. For example, one concern is that, despite their proficiency in producing functional code, LLMs often lack awareness and understanding of secure coding practices [10][11]. Another concern is that LLMs are trained on enormous datasets, often gathered from unverified online platforms (e.g., GitHub [12] and HuggingFace [13]). This reliance on potentially insecure data introduces data poisoning risks, where an attacker inserts harmful data samples into the training set [14]. These and other shortcomings can lead LLMs to generate insecure code, as well as influence their abilities to detect and fix security vulnerabilities.
@@ I Introduction
I-A Outline
The remainder of this paper is organized as follows. Section II describes related works. Section III describes the methodology we have followed in this work, as well as the research questions that guided our investigation. Section IV discusses the vulnerabilities identified in LLM-generated code, by the state of the art. Section V examines LLMs ability to detect code vulnerabilities, Section VI their ability to fix pre-identified vulnerabilities, and Section VII their potential to detect and fix vulnerabilities in a unified process. Furthermore, Section VIII covers prompting techniques used for vulnerability detection and fixing, as described by the studies covered in this review. Section IX explores the impact of poisoned training data on LLMs ability to generate secure code, detect vulnerabilities, and fix them. Section X addresses the research questions, open challenges, and threats to validity. Finally, Section XI concludes the paper.
II Related Work
In recent years, several literature reviews have focused on different aspects of the intersection between LLMs and code security. As we discussed in Section I, in this SLR we focus on three main thematic areas. Although other surveys have covered these areas, all of them focused on only one (or two) areas at a time. To the best of our knowledge, this is the first comprehensive survey that addresses all three areas. Furthermore, this is the first paper that attempts to categorize the different vulnerabilities that could be introduced by LLMs. For example, Negri et al. [15] conducted an SLR to explore the impact of AI models on software security, finding that LLMs often generate code containing specific vulnerabilities, but they did not provide a systematic categorization. Extending their work, we organize the vulnerabilities that can be introduced by LLM-generated code, into ten distinct categories. Besides, we analyze quantitatively their prevalence across studies, highlighting less explored areas.
@@ II Related Work
Yao et al. [16] reviewed various applications of LLMs in security, including vulnerability detection and fixing. However, their work did not investigate the impact of prompting techniques on such tasks, nor the merging of vulnerability detection and fixing in a unified process. Zhou et al. [17] provided an overview of prompting techniques, listing which papers applied each technique for vulnerability detection and repair. Unlike their work, we focus on how each prompting technique impacts LLMs effectiveness in these tasks.
Furthermore, the literature does not cover the impact of data poisoning attacks on LLMs code generation, and our work fills this gap by analyzing this potential impact. Even though data poisoning attacks are explored by Yao et al. [16], the paper does not discuss the effects of these attacks on LLMs secure code generation, nor on vulnerability-related tasks. Chen et al. [18] reviewed defenses and attacks targeting LLMs used for code generation, highlighting the general impact of such attacks. Although they also discussed the consequences of poisoning attacks on secure code generation, this was not the main focus of their paper. For example, they also explored the effects on tasks unrelated to security, including code summarization and code search. In contrast, we provide an in-depth exploration of security-specific aspects, discussing how poisoning attacks influence the ability of LLMs to generate secure code, detect vulnerabilities, and propose fixes.
A few relevant studies conducted broader literature reviews, that extend beyond the focus on code security. Xu et al. [19] examined LLM applications across multiple cybersecurity tasks, including vulnerability detection, malware analysis, network intrusion detection, and phishing detection. Hou et al. [20] conducted a comprehensive study on the use of LLMs in software engineering, including, but not limited to, code security. They pointed out that code generation and program repair are the most common tasks where LLMs are applied to, in software development and maintenance. Last, Zhang et al. [21] focused on methodologies for constructing LLMs, specifically designed for cybersecurity tasks.
III Methodology
In this study, we conducted an SLR following the well-established guidelines proposed by Petersen et al. [22]. Our methodology, illustrated in Figure 1, includes three main phases: Planning, Conducting, and Reporting. In the remainder of this section, we detail the Planning phase.
Figure 1: Systematic literature review methodology.
III-A Research Questions
@@ III-B Search Strategy
III-C Selection Criteria
To ensure a relevant review of the literature, we defined specific inclusion and exclusion criteria. We designed the inclusion criteria to make sure the selected papers meet certain standards and relevance, while exclusion criteria help us eliminate papers that do not fit the research objectives or quality standards. The detailed criteria for paper selection are outlined in Table I.
TABLE I: Inclusion and exclusion criteria for paper selection.
@@ III-C Selection Criteria
IN2: Complete papers presented at specified conferences or published in journals, or available as preprints on arXiv.
@@
IV Security Vulnerabilities introduced by LLMs
Although LLMs became popular for their ability to generate functional code fast and speed up development, they can introduce security vulnerabilities that developers might miss, leading to risks regarding code integrity and safety [11]. In this section, we focus on specific vulnerabilities introduced by LLM-generated code, as identified through various studies we have analyzed in this paper.
In total, we have identified 20 papers that discuss vulnerabilities in LLM-generated code. We have collected the vulnerabilities highlighted in each paper, and organized them in 10 categories, shown in Table II. In this table, we indicate with a green checkmark () whether a vulnerability was identified in the study.
TABLE II: Vulnerabilities introduced by LLM generated code identified in the studies.
@@ IV Security Vulnerabilities introduced by LLMs
Liu et al. [23]
Khoury et al. [24]
Asare et al. [25]
Perry et al. [26]
Sandoval et al. [27]
Pearce et al. [28]
Elgedawy et al. [29]
Siddiq et al. [30]
Hajipour et al. [31]
Kim et al. [32]
He et al. [11]
Siddiq et al. [33]
Tóth et al. [34]
Tihanyi et al. [35]
Hamer et al. [36]
Rabbi et al. [37]
Fu et al. [38]
He et al. [39]
Mohsin et al. [10]
Gong et al. [40]
@@
Most of these studies identified vulnerabilities using the Common Weakness Enumeration (CWE) [41], with the corresponding CWE numbers listed in Table III. Many of these vulnerabilities are included in MITRE ’2023 CWE Top 25 Most Dangerous Software Weaknesses’ list [42], while others are referenced indirectly through related child CWEs.
TABLE III: Studies and the CWE Numbers.
@@
Liu et al. [23] 20, 22, 78, 79, 89, 119, 125, 190, 200, 206, 416, 434, 476, 502, 522, 787, 798
Khoury et al. [24] 35, 94, 133, 190, 256, 323, 338, 377, 400, 453, 502, 564, 687, 759
Asare et al. [25] 284, 664, 682, 691, 707, 710
Sandoval et al. [27] 252, 401, 416, 476, 758, 787
Pearce et al. [28] 20, 22, 77, 78, 79, 89, 119, 125, 190, 200, 306, 416, 434, 476, 522, 787, 798
Siddiq et al. [30] 79, 208, 209, 215, 287, 295, 312, 338, 798, 918
Hajipour et al. [31] 20, 22, 79, 190, 787
Kim et al. [32] 75, 119, 197, 401, 457, 467, 476, 562, 685, 686, 758, 788
He et al. [11] 22, 78, 79, 89, 125, 190, 476, 787
Siddiq et al. [33] 20, 22, 79, 89, 94, 297, 352, 327, 502, 601, 611, 730
Tóth et al. [34] 20, 80, 83
Tihanyi et al. [35] 120, 129, 190, 191, 476, 628, 761, 787, 788, 843
Hamer et al. [36] 327, 328, 404, 772, 798
Rabbi et al. [37] 20, 78, 94, 259, 327, 330, 400, 502, 605, 703
Fu et al. [38] 22, 78, 89, 94, 117, 215, 295, 312, 367, 369, 502, 563, 617, 628, 772, 798
He et al. [39] 22, 79, 89, 119, 200, 295, 326, 327, 338, 352, 502, 611, 732, 915
Mohsin et al. [10] 20, 22, 89, 200, 352, 676, 759
Gong et al. [40] 22, 78, 79, 80, 90, 94, 95, 99, 113, 116, 117, 200, 209, 215, 250, 252, 259, 269, 283, 285, 319, 321, 326, 329, 330, 331, 367, 377, 385, 400, 406, 414, 425, 434, 454, 462, 502, 522, 595, 601, 605, 611, 641, 643, 732, 759, 760, 776, 798, 827, 841, 918, 941, 943, 1204
Figure 2: Bar chart showing the number of studies discussing each category of security vulnerabilities introduced by LLMs.
IV-A Injection Vulnerabilities
@@
Two studies focused extensively on injection vulnerabilities. Khoury et al. [24] conducted an experiment where they generated 21 programs in 5 different programming languages using ChatGPT 3.5. The programs were intentionally diverse, each designed to highlight the risks associated with specific vulnerabilities. Among these, 6 programs specifically targeted injection vulnerabilities. Notably, in every instance, ChatGPT 3.5 failed to do proper input sanitization, resulting in various types of injection vulnerabilities. Similarly, Tóth et al. [34] focused on the security of 2500 PHP websites, generated by GPT-4. The research was conducted by analyzing the code from 2,500 PHP websites. Using both dynamic and static scanners, manual code verification, and penetration testing, they identified various vulnerabilities, particularly concerning the security of file upload functions, as well as SQL injections, XSS, and reflected XSS.
@@
The following CWEs related to injections were frequently identified throughout the studies: CWE-79: Improper Neutralization of Input During Web Page Generation (’Cross-site Scripting’) [23, 28, 33, 30, 31, 11, 39, 40], and CWE-89: Improper Neutralization of Special Elements used in an SQL Command (’SQL Injection’) [23, 28, 11, 33, 38, 39, 10].
IV-B Memory Management
@@
This category includes vulnerabilities related to memory management, identified across most of the included studies. These vulnerabilities include overflow vulnerabilities, such as buffer and integer overflow, mostly caused by the lack of input sanitization. Moreover, this category includes null pointer dereference and use-after-free memory allocation vulnerabilities. These issues arise due to improper handling of memory references and pose significant risks to system security. Although [23, 24, 26, 27, 28, 31, 32, 11, 33, 35, 39] all identified memory vulnerabilities in code produced by LLMs, Liu et al. [23] placed special importance on memory management vulnerabilities. They performed a systematic empirical assessment of the quality of code generated using ChatGPT, by solving 728 algorithmic problems in 5 languages. Their study highlighted significant issues in memory management within LLM-generated code. Specifically, they found that the majority of vulnerabilities were related to missing null pointer tests, with 91.8% of the total vulnerabilities attributed to the MissingNullTest query. These code snippets failed to check for null after memory allocation, which can lead to critical security vulnerabilities such as CWE-476: Null pointer dereference, also identified in [27, 28, 32, 35].
@@
In addition to null pointer dereference, other frequently identified CWEs related to memory management include: CWE-190: Integer Overflow or Wraparound [23, 24, 28, 31, 11, 35], and CWE-416: Use After Free [23, 27, 28].
IV-C File Management

### v3

# From Vulnerabilities to Remediation: A Systematic Literature Review of LLMs in Code Security

@@ Abstract.
$c$s: Security and privacy Software security engineeringccs: Security and privacy Vulnerability scannersccs: Computing methodologies Artificial intelligence
1. Introduction
Large Language Models (LLMs) are Machine Learning (ML) models designed for natural language processing (Blank, 2023). Representative models such as ChatGPT (Ope, [n. d.]), Llama (Touvron et al., 2023), GitHub Copilot (cop, 2024), and BERT (Kenton and Toutanova, 2019) have gained extreme popularity and recognition in the last couple of years due to their ability to perform well across different tasks.
Among other capabilities, LLMs effectively translate natural language into code, assist in code comprehension, and answer questions related to code (Kazemitabaar et al., 2024). Many developers utilize these abilities to optimize coding processes, reducing development time and thereby improving productivity. According to GitHub statistics, GitHub Copilot is now responsible for generating approximately 46% of code and boosts developers’ coding speed by up to 55% (Dohmke, 2024). In addition to their coding skills, LLMs can be powerful tools for vulnerability detection and fixing, particularly when provided with detailed prompts and clear background information (Le et al., 2024; Zhou et al., 2024d).
While LLMs represent an advancement in code generation, they exhibit several limitations that pose security risks. One concern is that, despite their proficiency in producing functional code, LLMs often lack awareness and understanding of secure coding practices (Mohsin et al., 2024; He and Vechev, 2023). Another concern is that LLMs are trained on enormous datasets, often gathered from unverified online platforms (e.g., GitHub (git, [n. d.]) and HuggingFace (hug, [n. d.])). This reliance on potentially insecure data introduces data poisoning risks, where an attacker inserts harmful data samples into the training set (Cotroneo et al., 2024). These and other shortcomings can lead LLMs to generate insecure code, as well as influence their abilities to detect and fix security vulnerabilities.
@@ 1. Introduction
1.1. Outline
This paper is organized as follows. Section 2 describes related works. Section 3 describes the methodology we have followed in this work. Section 4 discusses the vulnerabilities identified in LLM-generated code, by the state of the art. Section 5 examines LLMs ability to detect code vulnerabilities, Section 6 their ability to fix pre-identified vulnerabilities, and Section 7 their potential to detect and fix vulnerabilities in a unified process. Furthermore, Section 8 covers prompting techniques used for vulnerability detection and fixing, as described by the studies covered in this review. Section 9 explores the impact of poisoned training data on LLMs ability to generate secure code, detect vulnerabilities, and fix them. Section 10 addresses the research questions, open challenges, and threats to validity. Finally, Section 11 concludes the paper.
2. Related Work
While the Introduction outlined the key themes and goals of our survey, this section reviews related work and outlines how our survey builds on and extends previous efforts.
In recent years, several literature reviews have focused on different aspects of the intersection between LLMs and code security. As we discussed in Section 1, in this SLR we focus on three main thematic areas. Although other surveys have covered these areas, all of them focused on only one (or two) areas at a time. To the best of our knowledge, this is the first comprehensive survey that addresses all three areas. Furthermore, this is the first paper that attempts to categorize the different vulnerabilities that could be introduced by LLMs. For instance, Negri et al. (Negri-Ribalta et al., 2024) conducted an SLR to explore the impact of AI models on software security, finding that LLMs often generate code containing specific vulnerabilities, but they did not provide a systematic categorization. Extending their work, we organize the vulnerabilities introduced by LLM-generated code, into ten distinct categories. Besides, we analyze quantitatively their prevalence across studies, highlighting less explored areas.
@@ 2. Related Work
Yao et al. (Yao et al., 2024a) reviewed various applications of LLMs in security, including vulnerability detection and fixing. However, their work did not investigate the impact of prompting techniques on such tasks, nor the merging of vulnerability detection and fixing in a unified process. Zhou et al. (Zhou et al., 2024a) provided an overview of prompting techniques, listing which papers applied each technique for vulnerability detection and repair. Unlike their work, we focus on how each prompting technique impacts LLMs effectiveness in these tasks.
Furthermore, the literature does not cover the impact of data poisoning attacks on LLMs code generation, and our work fills this gap by analyzing this potential impact. Even though data poisoning attacks are explored by Yao et al. (Yao et al., 2024a), the paper does not discuss the effects of these attacks on LLMs secure code generation, nor on vulnerability-related tasks. Chen et al. (Chen et al., 2024b) reviewed defenses and attacks targeting LLMs used for code generation, highlighting the general impact of such attacks. Although they also discussed the consequences of poisoning attacks on secure code generation, this was not the main focus of their paper. For example, they also explored the effects on tasks unrelated to security, including code summarization and code search. In contrast, we provide an in-depth exploration of security-specific aspects, discussing how poisoning attacks influence the ability of LLMs to generate secure code, detect vulnerabilities, and propose fixes.
A few relevant studies conducted broader literature reviews extending beyond code security. Xu et al. (Xu et al., 2024) examined LLM applications across multiple cybersecurity tasks, including vulnerability detection, malware analysis, network intrusion detection, and phishing detection. Hou et al. (Hou et al., 2023) reviewed the use of LLMs in software engineering, including, but not limited to, code security. They noted that code generation and program repair are the most common applications in software development and maintenance. Finally, Zhang et al. (Zhang et al., 2024a) focused on methodologies for constructing LLMs, specifically designed for cybersecurity tasks.
3. Methodology
In this study, we conducted an SLR following the well-established guidelines proposed by Petersen et al. (Petersen et al., 2015). Our methodology, illustrated in Figure 1, includes three main phases: Planning, Conducting, and Reporting. In the remainder of this section, we detail the Planning phase.
Figure 1. Systematic literature review methodology.
3.1. Research Questions
@@ 3.2. Search Strategy
3.3. Selection Criteria
To ensure a relevant review of the literature, we defined specific inclusion and exclusion criteria. We designed the inclusion criteria to make sure the selected papers meet certain standards and relevance, while exclusion criteria help us eliminate papers that do not fit the research objectives or quality standards. The detailed criteria for paper selection are outlined in Table 1.
Table 1. Inclusion and exclusion criteria for paper selection.
@@ 3.3. Selection Criteria
IN2: Complete papers retrieved from searches in the IEEE Xplore Digital Library, ACM Digital Library, ScienceDirect, or SpringerLink as well as preprints available on arXiv.
@@
4. Security Vulnerabilities Introduced by LLMs
Although LLMs became popular for their ability to generate functional code fast and speed up development, they can introduce security vulnerabilities that developers might miss, leading to risks regarding code integrity and safety (He and Vechev, 2023). In this section, we focus on specific vulnerabilities introduced by LLM-generated code, as identified through various studies we have analyzed in this paper.
In total, we have identified 20 papers that discuss vulnerabilities in LLM-generated code. We have collected the vulnerabilities highlighted in each paper, and organized them in 10 categories, shown in Table 2. In this table, we indicate with a green checkmark () whether a vulnerability was identified in the study.
Table 2. Vulnerabilities introduced by LLM generated code identified in the studies.
@@ 4. Security Vulnerabilities Introduced by LLMs
Liu et al. (Liu et al., 2024c)
Khoury et al. (Khoury et al., 2023)
Asare et al. (Asare et al., 2023)
Perry et al. (Perry et al., 2023)
Sandoval et al. (Sandoval et al., 2023)
Pearce et al. (Pearce et al., 2022)
Elgedawy et al. (Elgedawy et al., 2024)
Siddiq et al. (Siddiq et al., 2024)
Hajipour et al. (Hajipour et al., 2024)
Kim et al. (Kim et al., 2024)
He et al. (He and Vechev, 2023)
Siddiq et al. (Siddiq and Santos, 2022)
Tóth et al. (Tóth et al., 2024)
Tihanyi et al. (Tihanyi et al., 2024)
Hamer et al. (Hamer et al., 2024)
Rabbi et al. (Rabbi et al., 2024)
Fu et al. (Fu et al., 2023)
He et al. (He et al., 2024)
Mohsin et al. (Mohsin et al., 2024)
Gong et al. (Gong et al., 2024)
@@
Figure 2. Bar chart showing the number of studies discussing each category of security vulnerabilities introduced by LLMs.
Most of these studies identified vulnerabilities using the Common Weakness Enumeration (CWE) (MITRE Corporation, [n. d.]), with the corresponding CWE numbers listed in Table 3. Many of these vulnerabilities are included in MITRE ’2023 CWE Top 25 Most Dangerous Software Weaknesses’ list (MITRE Corporation, 2023), while others are referenced indirectly through related child CWEs.
Table 3. Studies and the CWE Numbers.
@@
Liu et al. (Liu et al., 2024c) 20, 22, 78, 79, 89, 119, 125, 190, 200, 206, 416, 434, 476, 502, 522, 787, 798
Khoury et al. (Khoury et al., 2023) 35, 94, 133, 190, 256, 323, 338, 377, 400, 453, 502, 564, 687, 759
Asare et al. (Asare et al., 2023) 284, 664, 682, 691, 707, 710
Sandoval et al. (Sandoval et al., 2023) 252, 401, 416, 476, 758, 787
Pearce et al. (Pearce et al., 2022) 20, 22, 77, 78, 79, 89, 119, 125, 190, 200, 306, 416, 434, 476, 522, 787, 798
Siddiq et al. (Siddiq et al., 2024) 79, 208, 209, 215, 287, 295, 312, 338, 798, 918
Hajipour et al. (Hajipour et al., 2024) 20, 22, 79, 190, 787
Kim et al. (Kim et al., 2024) 75, 119, 197, 401, 457, 467, 476, 562, 685, 686, 758, 788
He et al. (He and Vechev, 2023) 22, 78, 79, 89, 125, 190, 476, 787
Siddiq et al. (Siddiq and Santos, 2022) 20, 22, 79, 89, 94, 297, 352, 327, 502, 601, 611, 730
Tóth et al. (Tóth et al., 2024) 20, 80, 83
Tihanyi et al. (Tihanyi et al., 2024) 120, 129, 190, 191, 476, 628, 761, 787, 788, 843
Hamer et al. (Hamer et al., 2024) 327, 328, 404, 772, 798
Rabbi et al. (Rabbi et al., 2024) 20, 78, 94, 259, 327, 330, 400, 502, 605, 703
Fu et al. (Fu et al., 2023) 22, 78, 89, 94, 117, 215, 295, 312, 367, 369, 502, 563, 617, 628, 772, 798
He et al. (He et al., 2024) 22, 79, 89, 119, 200, 295, 326, 327, 338, 352, 502, 611, 732, 915
Mohsin et al. (Mohsin et al., 2024) 20, 22, 89, 200, 352, 676, 759
Gong et al. (Gong et al., 2024) 22, 78, 79, 80, 90, 94, 95, 99, 113, 116, 117, 200, 209, 215, 250, 252, 259, 269, 283, 285, 319, 321, 326, 329, 330, 331, 367, 377, 385, 400, 406, 414, 425, 434, 454, 462, 502, 522, 595, 601, 605, 611, 641, 643, 732, 759, 760, 776, 798, 827, 841, 918, 941, 943, 1204
4.1. Injection Vulnerabilities
@@
Two studies focused extensively on injection vulnerabilities. Khoury et al. (Khoury et al., 2023) conducted an experiment where they generated 21 programs in 5 different programming languages using ChatGPT 3.5. The programs were intentionally diverse, each designed to highlight the risks associated with specific vulnerabilities. Among these, 6 programs specifically targeted injection vulnerabilities. Notably, in every instance, ChatGPT 3.5 failed to do proper input sanitization, resulting in various types of injection vulnerabilities. Similarly, Tóth et al. (Tóth et al., 2024) focused on the security of 2500 PHP websites, generated by GPT-4. Using both dynamic and static scanners, manual code verification, and penetration testing, they identified various vulnerabilities, particularly concerning the security of file upload functions, as well as SQL injections and XSS.
@@
The studies frequently identified the following injection-related CWEs: CWE-79: Improper Neutralization of Input During Web Page Generation (’Cross-site Scripting’) (Liu et al., 2024c; Pearce et al., 2022; Siddiq and Santos, 2022; Siddiq et al., 2024; Hajipour et al., 2024; He and Vechev, 2023; He et al., 2024; Gong et al., 2024), and CWE-89: Improper Neutralization of Special Elements used in an SQL Command (’SQL Injection’) (Liu et al., 2024c; Pearce et al., 2022; He and Vechev, 2023; Siddiq and Santos, 2022; Fu et al., 2023; He et al., 2024; Mohsin et al., 2024).
4.2. Memory Management
@@
This category includes vulnerabilities related to memory management, identified across most of the included studies. These vulnerabilities include overflow vulnerabilities, such as buffer and integer overflow, mostly caused by the lack of input sanitization. Moreover, this category includes null pointer dereference and use-after-free memory allocation vulnerabilities. These issues arise due to improper handling of memory references and pose significant risks to system security. Although (Liu et al., 2024c; Khoury et al., 2023; Perry et al., 2023; Sandoval et al., 2023; Pearce et al., 2022; Hajipour et al., 2024; Kim et al., 2024; He and Vechev, 2023; Siddiq and Santos, 2022; Tihanyi et al., 2024; He et al., 2024) all identified memory vulnerabilities in code produced by LLMs, Liu et al. (Liu et al., 2024c) placed special importance on memory management vulnerabilities. They performed a systematic empirical assessment of the quality of code generated using ChatGPT, by solving 728 algorithmic problems. Their study highlighted significant issues in memory management within LLM-generated code. Specifically, they found that the majority of vulnerabilities were related to missing null pointer tests, with 91.8% of the total vulnerabilities attributed to the MissingNullTest query. These code snippets failed to check for null after memory allocation, which can lead to critical security vulnerabilities such as CWE-476: Null pointer dereference, also identified in (Sandoval et al., 2023; Pearce et al., 2022; Kim et al., 2024; Tihanyi et al., 2024).
@@
In addition to null pointer dereference, other frequently identified CWEs related to memory management are: CWE-190: Integer Overflow or Wraparound (Liu et al., 2024c; Khoury et al., 2023; Pearce et al., 2022; Hajipour et al., 2024; He and Vechev, 2023; Tihanyi et al., 2024), and CWE-416: Use After Free (Liu et al., 2024c; Sandoval et al., 2023; Pearce et al., 2022).
4.3. File Management